OpenStudy (anonymous):

1 over 1+tanx + 1 over 1+cotx =1

OpenStudy (anonymous):

Verify the identities

OpenStudy (anonymous):

is it \[\frac{1}{1+\tan(x)}+\frac{1}{1+\cot(x)}=1\]?

OpenStudy (ybarrap):

Set tan(x) = a/b and cot(x) = b/a. Simplify. Result follows.

OpenStudy (anonymous):

Yea @satellite73

OpenStudy (dumbcow):

@Gabysolis49 cot = 1/tan \[\frac{1}{1+\tan x}+\frac{1}{1+\frac{1}{\tan x}}\] \[=\frac{1}{1+\tan x}+\frac{1}{\frac{\tan x}{\tan x}+\frac{1}{\tan x}}\] \[=\frac{1}{1+\tan x}+\frac{1}{\frac{\tan x +1}{\tan x}}\] \[=\frac{1}{1+\tan x}+\frac{\tan x}{1+\tan x}\] \[=\frac{1+\tan x}{1+\tan x}\] \[=1\]